OpenStudy (anonymous):

Find an equation for the line that is perpendicular to the line y = 2/7x + 3 and that passes through the point (3, −17).

OpenStudy (angela210793):

perpendicular---> m=-1/(2/7)=-7/2 y-y1=m(x-x1) y+17=-7/2(x-3)